OpenStudy (anonymous):

Use complete sentences to describe the steps taken to simplify this problem. Make sure you include information about the new common denominator, the final simplified expression, and any restrictions. 3/x - 3 /x+1 - 3 /x+ 2.

OpenStudy (anonymous):

\[\frac{ 3 }{ x } - \frac{ 3 }{ x-1 } - \frac{ 3 }{ x+2 }\]